# Data Integration with Fusion Searchlight: Classifying Brain States from Resting-state fMRI

### v1

@@ 3.1 Artificial dataset
We first investigated basic properties of the FuSL framework using the artificial dataset. Figure 2 (A) illustrates the 2 ROIs with artificial signals, whereby signal in ROI 1 is observed in sources 1 and 2, whereas the signal in ROI 2 is only observed in source 3, as depicted in figure 2 (B). We sampled signal amplitudes from a Gaussian normal distribution (with mean $\mu=0$ and standard deviation $\sigma=1$) and added noise with a power of $P_{noise}=1$, resulting in an average signal to noise ratio of $S/N=1$. An example of one sample with a noisy signal is shown in figure 2 (C). We then studied the decoding accuracy using different combinations of these sources as inputs for the FuSL. To fit the FuSL we used 10-fold stratified CV with 10 repetitions and used a SL radius of 3 vertices. As a decoder we employed a support vector classifier with a radial basis function kernel (Cortes and Vapnik, 1995, Schölkopf et al., 1998). Figure 2 (D) shows statistically significant decoding accuracy across the cortex when one incorporates all three sources simultaneously. In ROI 1 we observed higher decoding accuracy values than in ROI 2, because in ROI 1 the FuSL classifier could combine information simultaneously present in source 1 and 2. We then reconstructed the impact, as well as the feature-weighted impact of each source, as shown in figure 2 (E) and (F). Impact values of source 1 and 2 are higher in ROI 1 than in ROI 2, and of source 3 higher in ROI 2 than in ROI 1. Further we can see that feature-weighted impact maps reflected well the ground truth signals shown in figure 2 (B), whereby we observed some variability due to the added noise. Incorporating differences in features between groups in these maps also allowed us to reconstruct the change of the signal in each source. In this manner we were able to retrieve the increase of the signal in source 1 and decrease of the signal in source 2.
@@
Figure 2: Investigating basic properties of the FuSL framework using an artificial dataset. We induce artificial signals in two ROIs (A), whereby signals of source 1 and 2 overlap in ROI 1 and signals in source 3 are only present in ROI 2 (B). (C) We then generate samples by randomly varying the signal amplitude and adding Gaussian noise. (D) We observe higher decoding accuracy values in ROI 1 than in ROI 2. (E) Impact values of source 1 and 2 are higher in ROI 1 than ROI 2, and of source 3 higher in ROI 2 than ROI 1. (F) The feature-weighted impact displays increased values in source 1 and 3 and the decrease of values in source 2. (G) and (H) show decoding accuracies for all combinations of sources in ROI 1 and 2 respectively. (I) Decoding accuracies averaged across ROI 1 and 2. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies in figure (G), (H) and (I) are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
@@
We then studied the decoding accuracy under different combinations of informative and uninformative data sources as inputs for FuSL. For the statistical testing of accuracy differences across CV folds, we employed a paired t-test incorporating variance correction proposed by Nadeau and Bengio (1999) to account for correlations in the training data across folds, and repeated 10-fold CV 10 times to obtain a reliable estimate (Bouckaert and Frank, 2004). We applied false discovery rate (FDR) correction to account for multiple comparisons of different combinations of sources (Benjamini and Hochberg, 1995). Figure 2 (G) shows the average decoding accuracies in ROI 1 for all combinations of the three modalities. We defined a combination of two sources using the ||||| | operator, indicating a concatenation of their features. Using only source 1 or source 2 as input leads to an accuracy of $82.5\%$ and $81.1\%$ respectively, whereby using their combination 1||||| |2 in FuSL increases the accuracy to $90.6\%$ (Cohen’s $d=0.856$ and $d=0.729$ respectively). Adding data source 3, which has no signal in ROI 1 slightly decreases the decoding accuracy to $90.2\%$ in ROI 1 (Cohen’s $d=0.043$). In ROI 2 an average accuracy of $81.9\%$ is achieved when using only the informative source 3 as an input, as shown in figure 2 (H). The performance decreases slightly but significantly to $80.3\%$ and $80.2\%$ when noisy, uninformative sources are added (Cohen’s $d=0.138$ and $d=0.150$ respectively), and to $79.7\%$ when two uninformative sources are added (Cohen’s $d=0.200$). Figure 2 (I) shows the average decoding accuracy of ROI 1 and 2. We observed that a combination of all sources 1||||| |2||||| |3 achieved the overall best decoding accuracy. The increase in decoding accuracy by adding an informative source is thereby considerably larger than the decrease when adding an uninformative source, even in this scenario of relatively high noise levels ($S/N=1$). We observed a similar behavior when combining sources with higher ($S/N=2$) and lower ($S/N=0.1$) signal to noise ratio in supplement II figure S1.
@@
We further demonstrate in an example, how reconstructing feature importance using Shapley values increased also the spatial specificity of detected regions that were relevant for decoding a brain state. To this aim, we generated a signal in one vertex on the surface as shown in figure 3 (A). Current applications typically report the entire area of SLs with significant decoding accuracy (Etzel et al., 2013, Viswanathan et al., 2012), which is shown in our example in figure 3 (B). This area can be considerably larger than the area of the actual signal, because all SLs that contain the informative vertex may be able to successfully decode. Computing feature importance on vertex-level based on Shapley values allowed us to trace back the exact location of an important vertex within a SL. If we then average importance values across overlapping SLs, and weight these maps with the decoding accuracy, the map of feature impact can precisely identify the original location of the signal, as shown in figure 3 (C).
@@ 3.2 Resting-state fMRI dataset
Here, we study how integrating complementary rs-fMRI metrics in FuSL can improve the prediction of pharmacological treatment in rs-fMRI data. We incorporated combinations of regional homogeneity (ReHo), fractional amplitude of low frequency fluctuations (fALFF) and functional connectivity efficiency (FCE) with the objective to predict whether participants have received the sedative alprazolam or a placebo. We split the data of 34 subjects into train and test data using stratified 10-fold CV. For statistical comparison of model accuracies we repeated 10-fold CV 100 times and used a paired t-test including variance correction (Nadeau and Bengio, 1999) and FDR correction (Benjamini and Hochberg, 1995). Figure 4 (A) shows the performance of FuSL related to different combinations of rs-fMRI metrics. Decoding accuracies averaged across the cortex are depicted in figure 4 (A1). Respective p-values and effect sizes are provided in supplement table S1. We further evaluated the number of vertices where FuSL was able to find significant differences between groups, using 10-fold CV and TFCE permutation-based correction with 1000 permutations. Average TFCE weighted accuracy values and the proportion of significant vertices are shown in figure 4 (A2) and (A3) respectively. The combinations ReHo||||| |FCE and ReHo||||| |fALFF||||| |FCE achieved the highest average decoding accuracies, whereby threshold-free cluster enhanced accuracies and the number of significant (TFCE corrected) vertices were highest for the combination ReHo||||| |fALFF||||| |FCE.
@@
Figure 4: Decoding performance of FuSL in dependency of metric combinations. (A) Decoding performance averaged across the whole cortex. (A1) Combinations of fALFF||||| |FCE and ReHo||||| |fALFF||||| |FCE achieved the highest average decoding accuracies. (A2) The combination ReHo||||| |fALFF||||| |FCE yielded highest threshold-free cluster enhanced accuracy values. (A3) The combination ReHo||||| |fALFF||||| |FCE also produced the highest number of significant vertices. (B) Areas that can be additionally decoded when using the combination ReHo||||| |fALFF||||| |FCE in FuSL. Locations where only ReHo||||| |fALFF||||| |FCE yielded significant decoding accuracies are highlighted in red in the left (L) and right (R) cortex. Locations where ReHo||||| |fALFF||||| |FCE and the respective other combination can both significantly decode are marked in white. Locations where only the respective other combination is able to decode are depicted in blue. Significant differences of accuracies in figure (A) are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ns: not significant.
@@
To illustrate which significant vertices could be additionally detected when combining all three rs-fMRI metrics, figure 4 (B) shows the differences between ReHo||||| |fALFF||||| |FCE and all other possible combinations of metrics. When using ReHo, FCE or ReHo||||| |FCE, no area on the cortex showed significant decoding accuracy. By using a combination of all three rs-fMRI metrics (ReHo||||| |fALFF||||| |FCE) we could find significant alterations in various typical resting-state networks (Yeo et al., 2011), as highlighted in red in figures 4 (B1), (B2) and (B3). Due to the administration of alprazolam we found modulations in large parts of the visual resting-state network across the full occipital cortex. In the default mode network we observe effects in the parietal cortex, inferior frontal gyrus, superior temporal gyrus, and the anterior and posterior cingulate cortex. In the dorsal attention resting-state network we found differences within the inferior temporal cortex, superior parietal lobule, angular gyrus and left precentral gyrus. Further we observe modulations predominantly in the frontal parts of the frontoparietal network, and modulations of the ventral attention network mainly in the anterior and posterior cingulate gyrus. In the somatomotor resting-state network we found an effect in the precentral gyrus. A detailed overlay of regions with these significant decoding accuracies and the resting-state atlas defined by Yeo et al. (2011) is provided in supplement II figure S2. Figure 4 (B4) shows the comparison to the case when using only fALFF as input. Both variants could detect significant alterations within the default mode network in the posterior cingulate cortex (overlap highlighted in white), but only using fALFF did not allow us to decode in any other region. Further the difference to the combination of ReHo||||| |fALFF is illustrated in figure 4 (B5). A combination of ReHo||||| |fALFF enabled us to also successfully decode within the visual, frontoparietal, ventral/dorsal attention and default resting-state network, but failed to decode in the motor cortex and large parts of the default network within the cingulate cortex and superior frontal gyrus. A small number of vertices that became significant incorporating a combination of ReHo||||| |fALFF but not ReHo||||| |fALFF||||| |FCE are marked in blue. Finally differences between fALFF||||| |FCE and ReHo||||| |fALFF||||| |FCE are depicted in figure 4 (B6). Here we can observe that on the one side a combination fALFF||||| |FCE was able to additionally decode some more vertices within the left and right somatosensory network and left limbic network. On the other side a combination of ReHo||||| |fALFF||||| |FCE allowed to capture more vertices within the default network in the superior temporal sulcus, right frontoparietal and dorsal attention network. Overall a combination of all three rs-fMRI metrics (ReHo||||| |fALFF||||| |FCE) yielded the highest number of vertices with statistically significant decoding accuracies.
@@
We then studied in detail the impact of each metric in the FuSL that combines all three rs-fMRI metrics (ReHo||||| |fALFF||||| |FCE). Figure 5 (A) shows statistically significant decoding accuracies across the cortex. Within these regions we can observe highest discrimination capabilities in the visual network within the extrastriate cortex and in the default mode network within left angular gyrus, right anterior cingulate cortex and bilateral in posterior cingulate cortex. Analyzing the impact and feature-weighted impact of individual rs-fMRI metrics allowed us to relate the decoding performance to the individual input metrics. A comparison between the impact and feature-weighted impact shows that the magnitudes of these measures coincide qualitatively very well (supplementary figure S3), but analyzing the feature-weighted impact additionally allows one to identify, whether a metric is increase or decreased due to the alprazolam treatment. In figure 5 (B1) we could observe that for ReHo a decrease within the visual network in the right extrastriate cortex is most indicative for the administration of alprazolam. Further we observed a characteristic decrease of ReHo mainly in frontal parts of the frontoparietal resting-state network, and within the default mode network a decrease in the parietal and posterior cingulate cortex. Simultaneously we found an increase in ReHo in the somatomotor network, and within the default mode network an increase in the anterior cingulate cortex. In comparison to ReHo, for fALFF we observed generally higher feature-weighted impact values across the cortex, as illustrated in figure 5 (B2). A characteristic increase of fALLF due to alprazolam can be found within the visual network bilateral in the extrastriate cortex. Additionally we found an increase with high impact in fALLF within the frontoparietal and default mode network, in frontal cortex, and anterior and posterior cingulate cortex. Related to alprazolam we found predominantly a decrease of FCE as shown in figure 5 (B3). This decrease of FCE has highest impact on the decoding also within the visual network in the extrastriate cortex, and within the default mode network in the superior temporal gyrus. Additionally an overlay of the decoding accuracy map and feature-weighted impact maps with the resting-state network atlas established by Yeo et al. (2011) is provided in supplement II figure S4.
@@
Figure 5: (A) Combining ReHo, fALFF and FCE in FuSL yields highest decoding accuracies in the visual network and frontal part of the default mode network. (B) The feature-weighted impact illustrates the informativeness, importance and relative increase or decrease of these metric, related to the treatment of alprazolam. (B1) Illustrates the decrease of ReHo mainly in the visual and frontoparietal network, including frontal parts of the default mode network. Simultaneously we observe an increase in ReHo in the somatomotor network and anterior cingulate cortex. (B2) A characteristic increase in fALFF can be found mainly in the visual, frontoparietal and default mode network. (B3) A high feature-weighted impact related to a decrease of FCE is highest within the visual network in the extrastriate cortex and default mode network in the superior temporal gyrus.

### v2

@@ 3.1 Artificial dataset
We first investigated basic properties of the FuSL framework using the artificial dataset with a moderate signal to noise ratio of $S/N=1$, a signal variability of $\sigma=1$ across samples and with no spatial smoothing. Figure 2 (A) illustrates the 2 ROIs with artificial signals, whereby signal in ROI 1 was observed in sources 1 and 2, whereas the signal in ROI 2 was only observed in source 3, as depicted in figure 2 (B). An example of one sample with a noisy signal is shown in figure 2 (C). We then studied the decoding accuracy using different combinations of these sources as inputs for the FuSL. To fit the FuSL we used 10-fold stratified CV with 10 repetitions and used a SL radius of 3 vertices. As a decoder we employed a support vector classifier (SVC) with a radial basis function (RBF) kernel (Cortes and Vapnik, 1995, Schölkopf et al., 1998). Figure 2 (D) shows statistically significant decoding test accuracies across the cortex when one incorporates all three sources simultaneously. In ROI 1 we observed higher decoding accuracy values than in ROI 2, because in ROI 1 the FuSL classifier could combine information simultaneously present in source 1 and 2. We then reconstructed the impact, as well as the feature-weighted impact of each source, as shown in figure 2 (E) and (F). Impact values of source 1 and 2 were higher in ROI 1 than in ROI 2, and impact of source 3 was higher in ROI 2 than in ROI 1. Further we can see that feature-weighted impact maps reflected well the ground truth signals shown in figure 2 (B), whereby we observed some variability due to the added noise. Incorporating differences in features between groups in these maps also allowed us to reconstruct the change of the signal in each source. In this manner, we were able to retrieve the increase of the signal in source 1, and a decrease of the signal in source 2.
@@
Figure 2: Investigating basic properties of the FuSL framework using an artificial dataset. We induce artificial signals in two ROIs (A), whereby signals of source 1 and 2 overlap in ROI 1 and signals in source 3 are only present in ROI 2 (B). (C) We then generate samples by randomly varying the signal amplitude and adding Gaussian noise. (D) We observed higher decoding test accuracy values in ROI 1 than in ROI 2. (E) Impact values of source 1 and 2 are higher in ROI 1 than ROI 2, and of source 3 higher in ROI 2 than ROI 1. (F) The feature-weighted impact displays increased values in source 1 and 3 and the decrease of values in source 2. (G) and (H) show decoding accuracies for all combinations of sources in ROI 1 and 2 respectively. (I) Decoding accuracies averaged across ROI 1 and 2. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies in figure (G), (H) and (I) are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
@@
We then studied the decoding accuracy under different combinations of informative and uninformative data sources as inputs for FuSL. Figure 2 (G) shows the average test accuracies in ROI 1 for all combinations of the three modalities, using 10 repetitions of 10-fold CV. We defined a combination of two sources using the ||||| | operator, indicating a concatenation of their features. Using only source 1 or source 2 as input yielded an accuracy of $82.5\%$ and $81.1\%$ respectively, whereby using their combination 1||||| |2 in FuSL increased the accuracy to $90.6\%$ (Cohen’s $d=0.856$ and $d=0.729$ respectively). In ROI 2 an average accuracy of $81.9\%$ was achieved when using only the informative source 3 as an input, as shown in figure 2 (H). The performance decreased slightly but significantly to $80.3\%$ and $80.2\%$ when noisy, uninformative sources were added (Cohen’s $d=0.138$ and $d=0.150$ respectively), and to $79.7\%$ when two uninformative sources were added (Cohen’s $d=0.200$). Figure 2 (I) shows the average decoding accuracy of ROI 1 and 2. We observed that a combination of all sources 1||||| |2||||| |3 achieved the overall best decoding accuracy. The increase in decoding accuracy by adding an informative source was thereby considerably larger than the decrease when adding an uninformative source. Figure 3 shows the statistical power of the decoding in ROI 1 in dependency of the sample size. Adding the informative source 2 to source 1 considerably increased the statistical power, especially when sample sizes were smaller. Adding an uninformative source 3 to source 1 again led to a slight decrease in statistical power.
In addition we compared the performance of the SVC to different MVPA decoding models such as a k-nearest neighbors classifier, a random forest classifier and ridge regression classifier in figure S2 (A). We also studied the impact of different kernel functions for SVC, and we found that a SVC with RBF kernel overall yielded the best results. A comparison of SL radii in figure S2 (B) revealed that a neighborhood order of 3 provided a good tradeoff between decoding accuracy and computational time. A more comprehensive evaluation of the impact of data fusion across datasets with different signal to noise ratios ($S/N=[1,0.1,0.05]$), different amounts of spatial autocorrelation (smoothing with $FWHM=[2mm,4mm,6mm]$), and different signal standard deviations across samples ($\sigma=[1,2]$) is additionally provided in figure S3 and S4. Spatial smoothing showed to especially improve decoding accuracy when noise levels are higher, but the overall characteristic accuracy gain when combining informative data sources was preserved across all datasets, demonstrating a good reliability of FuSL across different scenarios.
Figure 3: Statistical power of the decoding analysis in ROI 1. Adding the informative source 2 to source 1 increased the power, especially when the sample size is smaller, whereas adding the uninformative source 3 led to a decrease in statistical power. Errorbars represent 95% confidence intervals across dataset subsampling repetitions.
@@
We further show in an example, how reconstructing feature importance using Shapley values increased also the spatial specificity of detected regions that were relevant for decoding a brain state. To this aim, we generated a signal in one vertex on the surface as shown in figure 4 (A). Current applications typically report the entire area of SLs with significant decoding accuracy (Etzel et al., 2013, Viswanathan et al., 2012), which is shown in our example in figure 4 (B). Such an area can be considerably larger than the area of the actual signal, because all SLs that contain the informative vertex may be able to successfully decode. Computing feature importance on vertex-level based on Shapley values allowed us to trace back the exact location of an important vertex within a SL. After averaging importance values across overlapping SLs, and weighting these maps with the decoding accuracy, the map of feature impact precisely identified the original location of the signal, as shown in figure 4 (C).
@@ 3.2 Resting-state fMRI dataset
In this application we studied how integrating complementary rs-fMRI metrics in FuSL could improve the prediction of pharmacological treatment in rs-fMRI data. We incorporated combinations of regional homogeneity (ReHo), fractional amplitude of low frequency fluctuations (fALFF) and functional connectivity efficiency (FCE) with the objective to predict whether participants have received the sedative alprazolam or a placebo. Figure 5 shows the performance of FuSL related to different combinations of rs-fMRI metrics. The decoding test accuracies averaged across the entire cortex and across 100 repetitions of 10-fold CV are depicted in figure 5 (A). Statistically significant differences of the average test accuracies are highlighted in figure 5 (A) and the respective p-values and effect sizes are provided in supplementary table S1. We further evaluated the number of vertices in which FuSL was able to find significant differences between groups, using 10-fold CV. The average TFCE weighted test accuracy values and the proportion of significant vertices are shown in figure 5 (B) and (C) respectively. We found that mainly combining the information in fALFF with a connectivity based rs-fMRI metrics (ReHo, FCE or both) improved decoding accuracies considerably. Overall highest decoding accuracy was achieved by a the combination ReHo||||| |fALFF||||| |FCE and the highest number of vertices with statistical significant decoding accuracy by the combination ReHo||||| |fALFF.
@@
Figure 5: Decoding performance of FuSL in dependency of rs-fMRI metric combinations. The combinations fALFF||||| |FCE, fALFF||||| |ReHo and ReHo||||| |fALFF||||| |FCE achieved highest average decoding test accuracies (A), highest threshold-free cluster enhanced accuracy (B) and largest number of significant vertices (C). (D) Local differences in decoding test accuracies between ReHo||||| |fALFF||||| |FCE and all other combinations in the left (L) and right (R) cortex. (E) Locations where only ReHo||||| |fALFF||||| |FCE yielded significant decoding accuracies are highlighted in red. Locations where only the respective other combination is able to decode are depicted in blue. Locations where ReHo||||| |fALFF||||| |FCE and the respective other combination can both significantly decode are marked in white. Significant differences of accuracies in figure (A) are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ns: not significant.
@@
To illustrate the effect of combining rs-fMRI metrics more specifically across the cortex, figures 5 (D) and (E) illustrate differences in decoding accuracies and statistically significant areas between ReHo||||| |fALFF||||| |FCE and all other combinations. In comparison to when incorporating FCE only, a combination of all three rs-fMRI metrics (ReHo||||| |fALFF||||| |FCE) could considerably increase decoding accuracy across the cortex and yielded significant alterations in various typical resting-state networks (Yeo et al., 2011), as shown in figures 5 (D1) and (E1). Due to the administration of alprazolam we found significant modulations in large parts of the visual resting-state network across the full occipital cortex (illustrated in red in figure 5 (E1)). In the default mode network we observed effects in the parietal cortex, inferior frontal gyrus, superior temporal gyrus, and the anterior and posterior cingulate cortex. In the dorsal attention resting-state network we found differences within the inferior temporal cortex, superior parietal lobule, angular gyrus and precentral gyrus. Further, we observed modulations predominantly in the frontal parts of the frontoparietal network, and modulations of the ventral attention network mainly in the anterior and posterior cingulate gyrus. In the somatomotor resting-state network we found an effect in the precentral gyrus. A detailed overlay of these comparisons with the resting-state atlas defined by Yeo et al. (2011) is provided in supplementary figure S5. Figures 5 (D2) and (E2) show the comparison to the case when using only ReHo for decoding. Using ReHo also allowed us to predict alprazolam treatment in the visual, default and parts of the right dorsal attention and motor network (overlap highlighted in white), but not in any other area. A small number of vertices mainly within the left ventral attention network showed significant decoding accuracy when incorporating ReHo only but not when using the combination ReHo||||| |fALFF||||| |FCE (highlighted in blue). A comparison to fALFF in figures 5 (D3) and (E3) highlights that both variants were able to detect significant alterations within the visual network in both hemispheres, but using only fALFF did not allow us to decode in large parts of the right default, motor, frontoparietal and dorsal attention network. As shown in figures 5 (D4) and (E4) adding FCE to ReHo made it possible to additionally predict the treatment in the superior temporal cortex, but reduced the prediction performance in the right frontal cortex. Finally, the comparison of ReHo||||| |fALFF||||| |FCE to the combinations fALFF||||| |FCE (figures 5 (D5) and (E5)) and ReHo||||| |fALFF (figures 5 (D6) and (E6)) illustrates that combinations of fALFF with a connectivity based metric, either ReHo or FCE or both, yielded overall very similar predictions, whereby ReHo added some information on modulations in the right default mode network, and FCE on modulations in the superior temporal cortex.
In addition we compared the SVC with RBF kernel to alternative MVPA decoding models, and also studied the impact of different SL raidii in supplementary figures S6 (A) and (B). We found that also on this rs-fMRI dataset a SVC with RBF kernel yielded best results, and a SL radius of order 3 provided a reasonable trade-off between performance and computation time. A comparison to alternative variants of the three rs-fMRI metrics illustrates in supplementary figure S6 (C) that fALLF was slightly more informative than ALFF, FC efficiency was slightly more informative than FC betweeness centrality, and ReHo with a neighborhood order of 5 was more characteristic than of order 3 for the alprazolam treatment, whereby these differences were not statistically significant. The effect of removing the within subject mean of each rs-fMRI metric is depicted in supplementary figure S7. Thereby removing this baseline variability of rs-fMRI metrics across subjects has shown to considerably improve the prediction of pharmacological treatment.
@@
We then studied in detail the impact of each metric in the FuSL that combined all three rs-fMRI metrics (ReHo||||| |fALFF||||| |FCE). Figure 6 (A) shows statistically significant decoding test accuracies across the cortex. Within these regions we observed highest discrimination capabilities in the visual network within the extrastriate cortex, in the default mode network within left angular gyrus, and within the right motor cortex. Additional performance metrics like decoding sensitivity, specificity and area under receiver operating characteristic (ROC) curve are visualized in supplementary figure S8. Analyzing the impact and feature-weighted impact of individual rs-fMRI metrics allowed us to relate the decoding performance to the individual input metrics. We therefore first compared two different variants for approximating Shapley values in supplementary figure S9, one based on permutation testing and one using linear explanation models (Lundberg and Lee, 2017). Both approaches yielded similar impact maps across metrics, whereby the permutation based method was considerably more computationally efficient. A comparison between the impact and feature-weighted impact shows that the magnitudes of these measures coincide qualitatively very well (supplementary figure S10), but analyzing the feature-weighted impact additionally allowed us to identify whether a metric increased or decreased under alprazolam treatment. In figure 6 (B1) we observed that for ReHo a decrease within the visual network was indicative of the administration of alprazolam. Further, we observed a characteristic decrease in ReHo mainly in the frontoparietal resting-state network, and within the default mode network a decrease in the parietal and posterior cingulate cortex. Simultaneously we found an increase in ReHo in the somatomotor network, and within the default mode network an increase in the anterior cingulate cortex and superior temporal cortex. Compared to ReHo, for fALFF we generally observed higher feature-weighted impact values across the cortex, as illustrated in figure 6 (B2). A characteristic increase in fALLF due to alprazolam was found bilaterally within the visual network in the extrastriate cortex. Additionally, we found an increase with high impact in fALLF within the frontoparietal and default mode network, in frontal cortex, and anterior and posterior cingulate cortex. In relation to alprazolam we found predominantly a decrease in FCE as shown in figure 6 (B3). This decrease in FCE had highest impact on decoding also within the visual network in the extrastriate cortex, and within the default mode network in the superior temporal gyrus. Additionally, an overlay of the decoding accuracy map and feature-weighted impact maps with the resting-state network atlas established by Yeo et al. (2011) is provided in supplementary figure S11.
@@
Figure 6: (A) Combining ReHo, fALFF and FCE in FuSL yielded highest decoding accuracies in the visual, default mode and motor network. (B) The feature-weighted impact illustrates the informativeness, importance and relative increase or decrease of these metric, related to the treatment of alprazolam. (B1) Illustrates the decrease of ReHo mainly in the visual and frontoparietal network, including frontal parts of the default mode network. Simultaneously we observe an increase in ReHo in the somatomotor network, superior temporal cortex and anterior cingulate cortex. (B2) A characteristic increase in fALFF was found widespread in the visual, frontoparietal and default mode network. (B3) A high feature-weighted impact related to a decrease of FCE was highest within the visual network and default mode network within the superior temporal gyrus.